OpenStudy (anonymous):

find two series solution of y''-2y'+y=0

OpenStudy (tkhunny):

Have you considered building the generalized series?

OpenStudy (abb0t):

You can solve this without using series to solve...must you use series?

OpenStudy (loser66):

look for \[y = \sum_{n=0}^\infty a_nx^n\] \[y'= \sum_{n=0}^\infty na_nx^{n-1}\] \[y"=\sum_{n=0}^\infty n(n-1)a_nx^{n-2}\] so, \[yours =\sum_{n=0}^\infty n(n-1)a_nx^{n-2}-2\sum_{n=0}^\infty na_nx^{n-1}+\sum_{n=0}^\infty a_nx^n\] now make them have the same exponent . Pay attention: the first term, exponent of x is n-2, you want to get n , that means you +2 into n (everywhere you see n, +2), The second term +1, the last term, keep it as it is. so you have \[\sum_{n=0}^\infty (n+2)(n+1)a_{n+2}x^{n}-2\sum_{n=0}^\infty (n+1)a_{n+1}x^{n}+\sum_{n=0}^\infty a_nx^n\] now you can put everything under the one sum \[\sum_{n=0}^\infty (n+2)(n+1)a_{n+2}x^{n}-2(n+1)a_{n+1}x^{n}+a_nx^n\] can you step up from here?

OpenStudy (loser66):

just factor x^n out, and let the coefficient =0, you can construct the relationship between \(a_{n+2}, a_{n+1}~~and~~~a_n\)

OpenStudy (loser66):

yes, but it's part of the program. Cannot say no to it

OpenStudy (abb0t):

What program?!

OpenStudy (loser66):

ODE

OpenStudy (abb0t):

Your solution is: y(x) = c\(_1\)e\(^x\)+c\(_2\)e\(^x\)x

OpenStudy (abb0t):

It should match something very similar to that. Best of luck! Cheers.

OpenStudy (loser66):

hehehe... fortunately, it's not mine. and it's not that.

OpenStudy (abb0t):

Your series solution will look different, but it should match the series to e\(^x\) and e\(^x\) x

OpenStudy (goformit100):

@priyeshvs "Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

\[y(x) = c_1 e^x+c_2 e^x x\]

OpenStudy (anonymous):

Solving by series-solution-method will give you something like \[y=a_0\left(1+x+\frac{1}{2!}x^2+\frac{1}{3!}x^3+\cdots\right)+a_1\left(x+x^2+\frac{1}{2!}x^3+\frac{1}{3!}x^4+\cdots\right)\] which is the same as \[y=C1e^x+C_2xe^x\] where \(a_0=C_1\) and \(a_1=C_2\).